OpenStudy (anonymous):

One number is 11 more than twice another number. If the sum of the numbers is twice their difference, find the numbers.

OpenStudy (academicgurusinc):

Let's say that the first number is A, and the second one is B A is 11 more than twice of B A = 11 + 2B ... (1) . The sum of the numbers (A + B) is twice their difference (A - B) A + B = 2(A-B) A + B = 2A - 2B A - 3B = 0 ... (2) . Substituting (1) into (2): 11 + 2B - 3B = 0 11 - B = 0 -> B = 11 . Then, insert the value of B into either (1) or (2) (I'm using (1) here): A = 11 + 2B = 11 + 2*11 = 11 + 22 A = 33 Hope that helps. If you found this helpful, I encourage to subscribe to our youtube channel ( https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg), to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)